OpenStudy (anonymous):

2(2x - 4) = 3(2x - 4)

OpenStudy (anonymous):

If what u say above is true then Mathemathics is no more..cuz the 2=3..so 1=0...so 1=2=0...so basically all numbers equal 0! :D

OpenStudy (smokeysthename):

do you have to solve for x?

OpenStudy (smokeysthename):

@tanker

OpenStudy (anonymous):

yes smokey

OpenStudy (smokeysthename):

okay, the first step is to distribute the 2 into 2x-4 and the 3 into 2x-4 do that and tell me what you get..

OpenStudy (anonymous):

4x-8 and 6x-12

OpenStudy (smokeysthename):

correct (: so you have 4x-8=6x-12 you have to get the x's to one side so by doing that you can either add 12 to each side or add 8 to each side

OpenStudy (smokeysthename):

which would you like to do?

OpenStudy (smokeysthename):

really hero? i got this.

OpenStudy (anonymous):

that wasnt right any way bo

OpenStudy (smokeysthename):

anyways, i added 12 to both sides so you'll get 4x+4=6x understand so far?

OpenStudy (smokeysthename):

^^how can he understand them if you just blow through them like that?

OpenStudy (anonymous):

new question 5[12 - (6 - 3a) - 2a] = 2(1 - a)

OpenStudy (smokeysthename):

start a new question and ill help you if you want me to

OpenStudy (anonymous):

5[12 - (6 - 3a) - 2a] = 2(1 - a)

OpenStudy (smokeysthename):

close this question and start a new one, then tag me